2.033g of a soluble salt is added to 51.6g of water in a calorimeter. The temperature changes from 27.9 °C to 23.2 °C. Calculate the heat change q of the reservoir in joules. Assume the specific heat capacity of the reservoir to be 4.184J/g. Hint: Answer includes no decimal places. Consider the sign of q.
Answer #1

q = m x C x ∆T

q = heat absorbed = ?

m = mass = 2.033 g + 51.6 g = 53.633 g

C = specific heat = 4.184 J/g/deg

∆T = change in temperature = 27.9 - 23.2 = 4.7 degrees (temperature decreases meaning an endothermic reaction)

q = ( 53.633 g)(4.184 J/g/deg)(4.7 deg) = 1054.6822184 J (positive sign since reaction is endothermic)

Answer: q = 1055 J

Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")